OpenStudy (anonymous):

a shoppper in a supermarket pushes a loaded 32kg cart with a horizontal force of 12N how far will it move in 3.5s starting from rest?

OpenStudy (anonymous):

\[F = ma\] \[a = \frac{ F }{ m } = 12/35 m/s^2\] \[s = ut + at^2/2\] here u = 0 \[S = at^2\]

OpenStudy (anonymous):

nw just subs the value and find s

OpenStudy (anonymous):

what does S stand for?

OpenStudy (anonymous):

displacement. that is what you were asked to find